40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Entry  Mon Aug 10 02:11:47 2015, Koji, Electronics, AM Stabilized EOM Driver, RF AM Measurement Unit E1500151 IMG_20150803_223403975.jpgIMG_20150803_221210816_HDR.jpgIMG_20150803_223420267.jpg
    Reply  Mon Aug 10 11:39:40 2015, Koji, Electronics, AM Stabilized EOM Driver, RF AM Measurement Unit E1500151 
       Reply  Mon Aug 10 11:57:17 2015, Koji, Electronics, AM Stabilized EOM Driver, RF AM Measurement Unit E1500151 screen_shot.png
          Reply  Mon Aug 10 12:09:49 2015, Koji, Electronics, AM Stabilized EOM Driver, RF AM Measurement Unit E1500151 IMG_20150809_215628585_HDR.jpg
             Reply  Fri Aug 28 01:08:14 2015, Koji, Electronics, AM Stabilized EOM Driver, RF AM Measurement Unit E1500151 ~ Calibration RFAM_detector_calib.pdfRFAM_detector_calib_spectra.pdf
                Reply  Fri Aug 28 02:14:53 2015, Koji, Electronics, AM Stabilized EOM Driver, RF AM Measurement Unit E1500151 ~ 37MHz OCXO AM measurement OCXO_AM_noise.pdf
                   Reply  Tue Sep 8 10:55:31 2015, Koji, Electronics, AM Stabilized EOM Driver, RF AM Measurement Unit E1500151 ~ 37MHz OCXO AM measurement 
Message ID: 231     Entry time: Mon Aug 10 02:11:47 2015     Reply to this: 232
Author: Koji 
Type: Electronics 
Category: AM Stabilized EOM Driver 
Subject: RF AM Measurement Unit E1500151 

This is an entry for the work on Aug 3rd.

LIGO DCC E1500151

Power supply check

- Removed the RF AM detector board and exposed the D0900848 power board. The board revision is Rev. A.

- The power supply voltage of +30.2V and -30.5V were connected as +/-31V supplies. These were the maximum I could produce with the bench power supply I had. +17.2V and -17.1V were supplied as +/-17V supplies.

- Voltage reference: The reference voltages were not +/-10V but +/-17V. The cause was tracked down to the voltage reference chip LT1021-10. It was found that the chip was mechanically destroyed (Attachment 1, the legs were cut by me) and unluckily producing +17V. The chip was removed from the board. Since I didn't have any spare LT1021-10, a 8pin DIP socket and an AD587 was used instead. Indeed AD587 has similar performance or even better in some aspects. This fixed the reference voltage.

- -5V supply: After the fix of the reference voltage, I still did not have correct the output voltage of -5V at TP12. It was found that the backpanel had some mechanical stress and caused a leg of the current boost transister cut and a peeling of the PCB pattern on the component lalyer (Attachment 2). I could find some spare of the transister at the 40m. The transister was replaced, and the pattern was fixed by a wire. This fixed the DC values of the power supply voltages. In fact, +/-24V pins had +/-23.7V. But this was as expected. (1+2.74k/2k)*10V = 23.7V .

- VREFP Oscillation: Similarly to the EOM/AOM driver power supply board (http://nodus.ligo.caltech.edu:8080/OMC_Lab/225), the buffer stage for the +10V has an oscillation at 762kHz with 400mVpp at VREFP. This was fixed by replacing C20 (100pF) with 1.2nF. The cap of 680pF was tried at first, but this was not enough to completely elliminate the oscilation.

- VREFN Oscillation: Then, similarly to the EOM/AOM driver power supply board (http://nodus.ligo.caltech.edu:8080/OMC_Lab/225), the amplifier and buffer stage for the -10V has an oscillation at 18MHz with 60mVpp at VREFN. This was fixed by soldering 100pF between pin6 and pin7 of U6.

- Voltage "OK" signal: Checked the voltage of pin5 of U1 and U4 (they are connected). Nominally the OK voltage had +2.78V. The OK voltage turned to "Low (~0V)" when:
The +31V were lowered below +27.5V.
The +31V were lowered below -25.2V.
The +17V were lowered below +15.2V.
The +17V were lowered below -15.4V.

- Current draw: The voltage and current supply on the bench top supplies are listed below

+30.2V 0.09A, -30.5V 0.08A, +17.2V 0.21A, -17.1V 0.10A

- Testpoint voltages:

TP12(-5V) -5.00V
TP11(-15) -14.99V
TP10(-24V) -23.69V
TP9(-10V) -9.99V
TP5(-17V) -17.15V
TP6(-31V) -30.69

TP2(+31V) +30.37V
TP3(+17V) +17.24V
TP8(+10V) +10.00V
TP16(+28V) +28.00V
TP13(+24V) +23.70V
TP14(+15V) +15.00V
TP15(+5V) +5.00V

 

Attachment 1: IMG_20150803_223403975.jpg  259 kB  Uploaded Mon Aug 10 03:51:20 2015  | Show | Hide all | Show all
Attachment 2: IMG_20150803_221210816_HDR.jpg  285 kB  Uploaded Mon Aug 10 12:09:04 2015  | Show | Hide all | Show all
Attachment 3: IMG_20150803_223420267.jpg  415 kB  Uploaded Mon Aug 10 12:09:12 2015  | Hide | Hide all | Show all
IMG_20150803_223420267.jpg
ELOG V3.1.3-